OpenStudy (anonymous):

Write the number that makes the statement True. 5.6 m = __________ mm Choose one answer. a. 0.56 b. 0.0056 c. 560 d. 5,600

ganeshie8 (ganeshie8):

hint : 1m = 1000 mm

OpenStudy (anonymous):

D if 1m = 1000m then 5.6 * 1000 = 5,600

OpenStudy (anonymous):

D! 1m=1000m so.... 1000*5=5000+600=5,600

OpenStudy (anonymous):

yeah! THE EXACT SAME THING i just said...